# Benchmarking Cross-Domain Audio-Visual Deception Detection

### v1

@@ 3.3 Attention-Mixer Fusion
TABLE I: The results of single-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M).
@@
TABLE II: The results of multi-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M), for different generalization strategies.
@@
4 Experiments
In this part, extensive experiments are conducted to benchmark the cross-domain performances on public deception detection datasets. In the following, we sequentially describe the benchmark datasets & metrics (Sec. 4.1), implementation details (Sec. 4.2), benchmarking results (Sec. 4.3 - 4.4) and fusion performances (Sec. 4.5).
4.1 Databases and Metrics
4.2 Implementation Details
Model Selection. Models for audio and visual modalities were selected to fit the data volume. For face frames, we adopted ResNet18 [31] and Gate Recurrent Unit (GRU) [40] models for facial feature extraction and temporal modeling, respectively. Two-layer multilayer perception (MLP) [41] models were used for AUs, gaze, and affect feature representation. For the audio-based Mel spectrogram, we used the ResNet18 [31] model for time-frequency feature representation. For audio waveforms, the Wave2Vec [42] model was applied for audio feature extraction.
TABLE III: The fusion results of single-to-single cross-domain generalization accuracy (%).
@@ 4.3 Cross-domain Testing with Unimodal Features
Single-to-Single Domain. Specifically, the models were trained on one dataset from one domain and tested on the other dataset from another domain. The experiments were conducted on the four public datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M). As shown in Tabel I, for visual modalities, we extracted the most adopted visual features including face frames, and behavior features such as AUs, gaze and affect. For audio modality, Mel spectrogram and waveform were extracted. We also applied several different backbone networks as audio and visual encoders. We can observe that R and B1 datasets generalized the best on B2, and M generalized the best on R. Note that the B2 dataset was not adopted as a source domain dataset as the original dataset had too much noise and we cleaned it only for testing. On average, we can observe that the best result was achieved by using visual (AU+gaze+affect) features.
@@
Multi-to-Single Domain. Here we fully evaluate the performance of multi-to-single cross-domain generalization using unimodal features. We conducted experiments for different domain sampling strategies. As shown in Table II, for domain-simultaneous training, the best generalization performance is achieved by training on the Bag of Lies and MU3D datasets and testing on the Real-life Trial dataset (69.16%), which was also the case for domain-by-domain training strategy (67.29%). For the domain-alternating strategy, the best result was observed when transferring to the Box of Lies dataset using the rest three datasets for training (65.85%). The results showed that the best generalization performances were obtained when transferring to the real-world dataset and the gameshow dataset by training on the lab-based datasets. This was because the lab-based datasets (Bag of Lies and MU3D) are relatively clean compared to the real-world dataset (Real-life Trial) and the gameshow dataset (Box of Lies). However, in the opposite case, the generalization performance degraded, for example, R&M to B1 and R&B1 to M. Different domain sampling strategies reached their best average performance on different input features and backbone networks. To be specific, for both domain-simultaneous and domain-by-domain strategies, models trained on visual (AU+gaze+affect) features reached their highest accuracies, which were 55.82% and 56.67%, respectively. Using the domain-alternating strategy, the best accuracy of 56.25% was achieved by training on visual (face frames) features. We can observe that models trained on visual modalities outperformed those trained on audio modalities across all the generalization strategies. This may be due to the rich deceptive cues captured by visual modalities in the publicly available datasets.
@@ 4.4 Domain-Simultaneous with Gradient Reversal Layer (GRL)
TABLE IV: The fusion results of multi-to-single cross-domain generalization accuracy (%) for different generalization strategies.
@@ 4.5 Cross-domain Testing with Multimodal Fusion
Here we present multimodal fusion results of cross-domain testing to evaluate models’ generalization capacities.
@@
Single-to-Single Domain with Fusion. In this section, we conducted experiments of single-to-single cross-domain testing. Two types of fusion positions were involved including score-level and feature-level fusion. For feature-level fusion, multiple fusion methods were adopted such as simple concatenation, SE-Concat [44], Cross-Atten [37], MLP-Mixer [19], and Attention-Mixer fusion (Ours). To be specific, simple concatenation refers to the concatenation of the extracted features before the input to the classifier. SE-concat stands for the SE attention applied to the concatenated features. Cross-Atten means the crossmodal attention among input features by using the attention mechanism from the Transformer. MLP-Mixer uses the method in  [19] on the extracted features. The modalities and input include three types, visual (face frames), visual (AU+gaze+affect), and audio (Mel spectrogram). It results in four combinations of either two or three types of inputs. As shown in Table IV, among each type of input combination, the proposed Attention-Mixer fusion (Atten-Mixer) achieved the best results. The rest fusion methods showed comparable results, which were less performed than Atten-Mixer.
@@
Multi-to-Single Domain with Fusion. We benchmarked multi-to-single cross-domain generalization with different fusion methods by three types of cross-testing strategies. In total, seven sub-experiments were conducted for different domain combinations. The results are shown in Tabel IV. On average, the best accuracies were 58.43%, 59.13%, and 60.78% on domain-simultaneous, domain-alternating, and domain-by-domain strategies, respectively. Among these, for both domain-simultaneous and domain-by-domain strategies, the best results were achieved by taking visual (face frames) and visual (AU+gaze+affect) features as input, while the best result for domain-alternating was achieved by using visual (face frames), visual (AU+gaze+affect) and audio (Mel spectrogram) features. Taking a close look at the average fusion results, for each type of modality input, Atten-Mixer achieved the best among the six fusion methods. showing the effectiveness of the proposed method. By using Atten-Mixer, in general, the average result showed slightly better when using the domain-by-domain strategy and taking visual (face frames) and visual (AU+gaze+affect) features as input. To sum up, the results showed that on the current publicly available datasets, visual features were better when it came to multi-to-single cross-domain generalizability. However, the performance differences were not a large gap, and there were also no significant differences by comparing two-to-one domain and three-to-one domain cross-testing performances.
@@
Figure 4: Ablation study for attention-mixer layers. The number of layers 4, 5, 6, and 7 are compared. The modality and inputs for A, B, C, and D are in line with those in Table III from the top to the bottom.
@@
Ablation Study for Attention-Mixer Fusion Module. We conducted an ablation study for the proposed attention-mixer fusion module with the changes in the number of attention-mixer layers. The experiments were conducted on single-to-single domain testing, where the average accuracies were compared. As shown in Fig. 4, the number of attention-mixer layers was set to 4, 5, 6, and 7. The modalities and inputs in Table III are compared, where “A” had the inputs of Visual (Face frame) + Visual(AU+Gaze+Affect), “B” had Visual (Face frames) + Audio(Mel spectrogram), “C” had Visual (AU+Gaze+Affect) + Audio (Mel spectrogram), and “D” had Visual (Face frame) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram). The results showed that models with 6 attention-mixer layers achieved the best average accuracies, followed by 7 attention-mixer layers.
@@ 4.6 Discussion
Ethical Consideration. Developing deception detection using AI should emphasize respecting privacy, minimizing psychological harm, preventing discrimination, promoting transparency, etc. Researchers should follow appropriate regulations to develop and deploy AI systems for deception detection. Potential misuses and negative impacts include invasion of privacy, discrimination, erosion of trust, etc. Mitigating these risks requires responsible practices from researchers and developers.

### v2

@@ 4.2 Implementation Details
TABLE I: The results of intra-domain testing accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), and MU3D (M).
Modality & Inputs Method R to R B1 to B1 M to M
Visual (Face frames) ResNet18 70.10 57.78 55.15
Visual (AU+Gaze+Affect) MLP 62.87 55.84 55.23
Audio (Mel spectrogram) ResNet18 75.78 55.69 56.45
Visual (Face frames) concat 59.09 54.46 56.30
+ Visual (AU+Gaze+Affect) Atten-Mixer (Ours) 63.05 55.45 57.59
Visual (Face frames) concat 68.75 54.74 56.19
+ Audio (Mel spectrogram) Atten-Mixer (Ours) 71.00 56.44 57.14
Visual (AU+Gaze+Affect) concat 68.88 56.25 53.27
+ Audio (Mel spectrogram) Atten-Mixer (Ours) 69.30 56.78 55.23
Visual (Face frames) concat 69.33 58.43 55.24
+ Visual (AU+Gaze+Affect) Atten-Mixer(Ours) 70.52 59.41 56.88
+ Audio (Mel spectrogram)
TABLE II: The results of single-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M).
@@
TABLE III: The results of multi-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M), for different generalization strategies.
@@
TABLE IV: The fusion results of single-to-single cross-domain generalization accuracy (%).
@@ 4.3 Cross-domain Testing with Unimodal Features
Intra-Domain Testing. Before performing cross-domain testing, we show the results of intra-domain testing on Real-life Trial (R), Bag of Lies (B1), and MU3D (M) in Table I. It is important to note that we did not use the Box of Lies (B2) dataset for training due to concerns about its quality. Specifically, we performed 5-fold intra-domain testing and reported the average accuracies. ResNet18 and MLP were used as baseline models, as they are widely used as encoders in audio-visual deception detection. As expected, the performance of intra-domain testing varies across datasets, reflecting the differences in their levels of difficulty.
@@
Single-to-Single Domain. Specifically, the models were trained on one dataset from one domain and tested on the other dataset from another domain. The experiments were conducted on the four public datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), and MU3D (M). As shown in Tabel II, for visual modalities, we extracted the most adopted visual features including face frames, and behavior features such as AUs, gaze and affect. For audio modality, Mel spectrogram and waveform were extracted. We also applied several different backbone networks as audio and visual encoders. We can observe that R and B1 datasets generalized the best on B2, and M generalized the best on R. Note that the B2 dataset was not adopted as a source domain dataset as the original dataset had too much noise and we cleaned it only for testing. On average, we can observe that the best result was achieved by using visual (AU+gaze+affect) features.
@@
Multi-to-Single Domain. Here we fully evaluate the performance of multi-to-single cross-domain generalization using unimodal features. We conducted experiments for different domain sampling strategies. As shown in Table III, for domain-simultaneous training, the best generalization performance is achieved by training on the Bag of Lies and MU3D datasets and testing on the Real-life Trial dataset (69.16%), which was also the case for domain-by-domain training strategy (67.29%). For the domain-alternating strategy, the best result was observed when transferring to the Box of Lies dataset using the rest three datasets for training (65.85%). The results showed that the best generalization performances were obtained when transferring to the real-world dataset and the gameshow dataset by training on the lab-based datasets. This was because the lab-based datasets (Bag of Lies and MU3D) are relatively clean compared to the real-world dataset (Real-life Trial) and the gameshow dataset (Box of Lies). However, in the opposite case, the generalization performance degraded, for example, R&M to B1 and R&B1 to M. Different domain sampling strategies reached their best average performance on different input features and backbone networks. To be specific, for both domain-simultaneous and domain-by-domain strategies, models trained on visual (AU+gaze+affect) features reached their highest accuracies, which were 55.82% and 56.67%, respectively. Using the domain-alternating strategy, the best accuracy of 56.25% was achieved by training on visual (face frames) features. We can observe that models trained on visual modalities outperformed those trained on audio modalities across all the generalization strategies. This may be due to the rich deceptive cues captured by visual modalities in the publicly available datasets.
Discussion: (1) Wave2Vec outperforms MFCC in audio feature extraction due to its ability to capture richer, more meaningful representations, as demonstrated across various sampling methods. Additionally, larger and more diverse datasets like B1 and M enhance generalization, leading to improved performance when models trained on these datasets are tested on smaller or less challenging datasets like R. The R dataset, composed of courtroom trial videos, tends to yield better results, while B2, a game show dataset, often produces unsatisfactory outcomes due to its lower video quality and data imbalance. Overall, visual features perform better than audio features and models using Wave2Vec features achieve better results compared to those trained on MFCC.
(2) While domain-simultaneous sampling is typically expected to enhance generalization by promoting domain-invariant learning, our average results show that domain-by-domain sampling slightly outperforms it in these audio-visual deception detection tasks. This may be due to its ability to capture domain-specific variations, which are beneficial in this context. However, the effectiveness of each sampling strategy depends on the features being used; for example, Wave2Vec features perform better with domain-simultaneous sampling, while MFCC features benefit more from domain-by-domain sampling. This suggests that the optimal sampling strategy may vary depending on the feature set’s predictive power and the specific dataset characteristics.
(3) Visual features, particularly behavioral features like AUs, gaze, and affective cues, outperform audio features in most of our experiments. Behavioral features are more predictive in the majority of testing pairs, as seen in Table II, and achieved the best average performance across sampling strategies in Table III. Their rich information content, capturing subtle signs of deception, may explain their superior performance over the face or audio features. Due to the lower dimension compared with face and audio features, behavioral features make them more effective when training on small datasets for detecting deception.
@@ 4.4 Domain-Simultaneous with Gradient Reversal Layer (GRL)
TABLE V: The fusion results of multi-to-single cross-domain generalization accuracy (%) for different generalization strategies.
@@ 4.5 Cross-domain Testing with Multimodal Fusion
Single-to-Single Domain with Fusion. In this section, we conducted experiments of single-to-single cross-domain testing. Two types of fusion positions were involved including score-level and feature-level fusion. For feature-level fusion, multiple fusion methods were adopted such as simple concatenation, SE-Concat [50], Cross-Atten [44], MLP-Mixer [21], and Attention-Mixer fusion (Ours). To be specific, simple concatenation refers to the concatenation of the extracted features before the input to the classifier. SE-concat stands for the SE attention applied to the concatenated features. Cross-Atten means the crossmodal attention among input features by using the attention mechanism from the Transformer. MLP-Mixer uses the method in  [21] on the extracted features. The modalities and input include three types, visual (face frames), visual (AU+gaze+affect), and audio (Mel spectrogram). It results in four combinations of either two or three types of inputs. As shown in Table IV, among all the fusion methods, the proposed Attention-Mixer fusion (Atten-Mixer) achieved the best results on average.
@@
Multi-to-Single Domain with Fusion. We benchmarked multi-to-single cross-domain generalization with different fusion methods by three types of cross-testing strategies. In total, seven sub-experiments were conducted for different domain combinations. The results are shown in Tabel V. On average, the best accuracies were 58.43%, 59.13%, and 60.78% on domain-simultaneous, domain-alternating, and domain-by-domain strategies, respectively. Among these, for both domain-simultaneous and domain-by-domain strategies, the best results were achieved by taking visual (face frames) and visual (AU+gaze+affect) features as input, while the best result for domain-alternating was achieved by using visual (face frames), visual (AU+gaze+affect) and audio (Mel spectrogram) features. Taking a close look at the average fusion results, for each type of modality input, Atten-Mixer achieved the best among the six fusion methods. showing the effectiveness of the proposed method. By using Atten-Mixer, in general, the average result showed slightly better when using the domain-by-domain strategy and taking visual (face frames) and visual (AU+gaze+affect) features as input. To sum up, the results showed that on the current publicly available datasets, visual features were better when it came to multi-to-single cross-domain generalizability. However, the performance differences were not a large gap, and there were also no significant differences by comparing two-to-one domain and three-to-one domain cross-testing performances.
@@
Figure 4: Ablation study for attention-mixer layers. The number of layers 4, 5, 6, and 7 are compared. The modality and inputs for A, B, C, and D are in line with those in Table IV from the top to the bottom.
@@
Ablation Study for Attention-Mixer Fusion Module. We conducted an ablation study for the proposed attention-mixer fusion module with the changes in the number of attention-mixer layers. The experiments were conducted on single-to-single domain testing, where the average accuracies were compared. As shown in Fig. 4, the number of attention-mixer layers was set to 4, 5, 6, and 7. The modalities and inputs in Table IV are compared, where “A” had the inputs of Visual (Face frame) + Visual(AU+Gaze+Affect), “B” had Visual (Face frames) + Audio(Mel spectrogram), “C” had Visual (AU+Gaze+Affect) + Audio (Mel spectrogram), and “D” had Visual (Face frame) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram). The results showed that models with 6 attention-mixer layers achieved the best average accuracies, followed by 7 attention-mixer layers.
TABLE VI: The fusion results of multi-to-single cross-domain generalization accuracy (%) for multimodal inter-domain gradient matching (MM-IDGM). The numbers in the brackets show the improved accuracies compared with Atten-Mixer.
Atten-Mixer(Ours) 53.23 57.43 63.37 52.19 57.01 63.37 62.38 58.43
IDGM 55.22 57.45 59.85 51.41 61.34 59.41 60.61 57.90
Visual (Face frames) + Visual (AU+Gaze+Affect) MM-IDGM 57.85 58.42 60.40 52.50 65.22 62.05 61.42 59.69 (+1.26)
Atten-Mixer(Ours) 55.69 64.36 56.44 53.75 58.88 59.41 58.42 58.14
IDGM 54.43 60.40 61.07 51.40 59.21 59.88 59.75 58.02
Visual (Face frames) + Audio (Mel spectrogram) MM-IDGM 53.54 59.41 62.38 51.56 59.07 62.38 59.50 58.26 (+0.12)
Atten-Mixer(Ours) 55.08 55.45 58.42 54.37 52.34 61.39 60.40 56.78
IDGM 55.49 54.92 58.36 50.31 53.15 60.35 59.21 55.97
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) MM-IDGM 56.05 56.03 59.37 52.41 53.40 62.42 60.98 57.23 (+0.45)
Atten-Mixer(Ours) 52.31 59.41 58.42 52.50 58.88 61.39 60.40 57.62
IDGM 52.45 57.02 60.42 54.25 58.42 61.79 60.03 57.77
Visual (Face frames)+ Visual (AU+Gaze+Affect)+ Audio (Mel spectrogram) MM-IDGM 52.77 57.43 61.39 55.57 59.46 62.58 60.60 58.54 (+0.92)
Performance of MM-IDGM We proposed MM-IDGM and conducted experiments for multi-to-single cross-domain generalization, with the results presented in Table VI. Both the IDGM and MM-IDGM algorithms are developed based on the domain simultaneous sampling method. Using the proposed Atten-Mixer fusion method as the baseline, we applied both the original IDGM and the proposed MM-IDGM. The results show that MM-IDGM consistently improves average accuracies across different modalities and inputs. However, cross-domain generalization in deception detection remains highly challenging, as some accuracies even decreased after applying IDGM. This may be due to the difficulty in optimizing the inner product between different domains. While MM-IDGM may not always perform optimally in certain cases, such as testing on the B2 dataset where dataset quality affects the results, it generally shows better performance in cross-domain generalization.
4.6 Network Details
Here, we present the detailed network architectures in Table VII, including encoders, the fusion module, and classifiers. This network architecture is utilized for experiments on multi-to-single cross-domain generalization with fusion.
TABLE VII: Architecture details for different modalities and fusion module.
Modality & Inputs: Visual (Face Frames)
Input Size: [B, 3, 32, 224, 224]
(For each frame) ResNet-18 [:-2] $\rightarrow$ [B, 32, 7, 7, 512]
Average Pooling $\rightarrow$ [B, 32, 512]
(For sequence) GRU Layer $\rightarrow$ [B, 32, 512]
Fully connected layer $\rightarrow$ [B, 2]
Modality & Inputs: Visual (AU+Gaze+Affect)
Input Size: [B, 50, 64]
(On d-50) Linear(50, 64) + LN + ReLU
Linear(64, 128) + LN + ReLU
Linear(128, 1) $\rightarrow$ [B*64, 1]
Reshape to [B, 64, 1]
(On d-64) Linear(64, 32) + LN + ReLU
Linear(32, 16) + LN + ReLU
Linear(16, 2) $\rightarrow$ [B, 2]
Modality & Inputs: Audio (Mel Spectrogram)
Input Size: [B, 3, 224, 224]
ResNet-18 [:12] $\rightarrow$ [B, 7, 7, 512]
Average Pooling $\rightarrow$ [B, 32, 512]
Reshape to [B, 32*512]
Linear(32*512, 12) $\rightarrow$ [B, 2]
Atten-Mixer Fusion
Input size: face frame [B, 32, 512],
behavioral [B, 64, 1],
and audio [B, 32, 512]
Projection Layer: face frame $\rightarrow$[B, 64, 1],
$\rightarrow$ [B, 64, 1],
and audio $\rightarrow$ [B, 64, 1]
Concatenation $\rightarrow$ [B, 64, 3]
Atten-Mixer Fusion: number of patches $\rightarrow$ 64,
channels $\rightarrow$ 3,
and output dim $\rightarrow$ 128
Output shape $\rightarrow$ [B, 128]
Classifier: Linear(128, 64)+ReLU+Linear(64, 2) $\rightarrow$ [B, 2]
@@ 4.7 Discussion
Ethical Consideration. Developing deception detection using AI should emphasize respecting privacy, minimizing psychological harm, preventing discrimination, etc. Potential misuse may cause negative impacts on society. If these systems are deployed without consent, they may violate individuals’ privacy rights by collecting and analyzing sensitive personal data, speech patterns, facial expressions, and body language without proper safeguards. Researchers should follow appropriate regulations to develop and deploy deception detection systems. They should collect and use data responsibly, ensuring informed consent and protecting personal information. The development of the models should be transparent and provide clear explanations for how decisions are made.